The directors of Thor Limited have appointed you as their financial consultant. They are considering new investment projects and need you to calculate the cost of capital for the company. The present capital structure is as follows; 2 000 000 ordinary shares with a par value of R1.00 per share. These shares are currently trading at R2.50 per share and the latest dividend paid is 40 cents. An average dividend growth of 9% is maintained. 1 500 000 8% R2.00 preference shares, with a market value of R1.80 per share. R10 000 000 non-distributable reserves R2 000 000 7% debentures due in 6 years time and the current yield to maturity is 10%, and R1 000 000 15% bank loan. Additional information: The company has a beta of 2.1 a risk free rate of 7% and a return of market of 16%. The company tax rate is 30%. Required 3.1 Calculate the weighted average cost of capital using Gordon growth model to calculate the cost of equity. (7) 3.2

1
Expert's answer
2022-03-21T00:18:49-0400

"CARM=7+1.2(16-7)=17.8"


"r1=\frac{0.4(1+0.09)}{2.5}+0.09=0.2644"


"d2=2\times0.08=0.16"


"r2=\frac{0.16(1+0.09)}{1.8}+0.09=0.1869"


"capital=2 000 000\times1+1 500 000\times2+2 000 000 +10 000 000+1 000 000=18 000 000"


"WACC=\frac{2 000 000}{18 000 000}\times26.44+\frac{3 000 000}{18 000 000}\times18.69+\frac{10 000 000}{18 000 000}\times17.8+\frac{2 000 000}{18 000 000}\times10(1-0.3)+\frac{1 000 000}{18 000 000}\times15(1-0.3)=17.30"